OpenStudy (anonymous):

find exact value of tan165

OpenStudy (aum):

tan(165) = tan(120 + 45) = ?

OpenStudy (aum):

\[ \tan(165) = \tan(120 + 45) = \frac{\tan(120)+\tan(45)}{1 - \tan(120)\tan(45)} \\ \tan(45) = 1 \\ \tan(60) = \sqrt{3} \\ \text{tan is negative in the second quadrant. Therefore, } \tan(120) = -\sqrt{3} \\ \tan(165) = \frac{-\sqrt{3}+1}{1 +\sqrt{3}} = \frac{-\sqrt{3}+1}{1 +\sqrt{3}} * \frac{1-\sqrt{3}}{1 -\sqrt{3}} = \\ \frac{1-2\sqrt{3}+3}{1 - 3} = \frac{4-2\sqrt{3}}{-2} = -2 + \sqrt{3} \]

OpenStudy (anonymous):

THANK YOU SO MUCH